An aircraft shell is fired vertically ward with a muzzle velocity at 100ms.calculate


a=the maximum height it can obtain


b=the time taken to teach the height


c=the instantaneous velocities at the end of 20sec and 50sec

1
Expert's answer
2021-11-08T16:49:21-0500

(a) Let's take the upwards as the positive direction. Then, we can find the maximum height from the kinematic equation:


"v^2=v_0^2+2gh,""0=v_0^2+2gh,""h=\dfrac{-v_0^2}{2g}=-\dfrac{(1000\ \dfrac{m}{s})^2}{2\times(-9.8\ \dfrac{m}{s^2})}=51\ km."

(b) We can find the time that shell takes to reach the maximum height from the kinematic equation:


"v=v_0+gt,""0=v_0+gt,""t=\dfrac{-v_0}{g}=\dfrac{-1000\ \dfrac{m}{s}}{-9.8\ \dfrac{m}{s^2}}=102\ s."

(c) We can find the instantaneous velocity at the end of "20\ s" from the kinematic equation:


"v=v_0+gt=1000\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\times20\ s=804\ \dfrac{m}{s}."

We can find the instantaneous velocity at the end of "50\ s" from the kinematic equation:


"v=v_0+gt=1000\ \dfrac{m}{s}+(-9.8\ \dfrac{m}{s^2})\times50\ s=510\ \dfrac{m}{s}."